OpenStudy (anonymous):

solve using the substitution method. 3+4y=-10 x=10-8y

OpenStudy (anonymous):

put x=4-6y in first eqn, 2(4-6y) +3y =12 8-12y +3y =12 8-9y=12 9y=-4 y= -4/9 we know, x=4-6y putting y=-4/9 we get, x= 4 - 6(-4/9) = 4 - 2(-4/3) = 4 + 8/3 = 20/3 Ans. x=20/3 y=-4/9

OpenStudy (anonymous):

is that an ordered pair

OpenStudy (anonymous):

was it checked

OpenStudy (hba):

@mizzkuntry To check put the values of x in the orignal question and check if L.H.S=R.H.S

OpenStudy (anonymous):

solve using the substitution method. 3+4y=-10 x=10-8y

OpenStudy (anonymous):

has to be simplified